### v1

Abstract
This work develops a new method for computing a finite quantum system’s continuum states and observables by applying a subspace projection (or reduced basis) method used in model order reduction studies to “discretize” the system’s continuous spectrum. The method extracts the continuum physics from solving Schrödinger equations with bound-state-like boundary conditions and emulates this extraction in the space of the input parameters. This parameter emulation can readily be adapted to emulate other continuum calculations as well, e.g., those based on complex energy or Lorentz integral transform methods. Here, I give an overview of the key aspects of the formalism and some informative findings from numerical experimentation with two- and three-body systems, which indicates the non-Hermitian quantum mechanics nature of the method. A potential connection with (near-)optimal rational approximation studied in Math literature is also discussed. Further details are provided in a separate paper.
Introduction.
This study concerns a finite quantum system, such as a nucleus, atom, or molecule, governed by the underlying Hamiltonian operator $H$. Of central interest is the matrix element of the resolvent operator between two spatially localized sources at total energy $E$:
$\mathcal{A}(E,\bm{\theta})\equiv\left\langle\tilde{S}(\bm{\theta})\left|\left[M(E,\bm{\theta})\right]^{-1}\right|S(\bm{\theta})\right\rangle\ ,$ (2)
with $M(E,\bm{\theta})\equiv E-H(\bm{\theta})$, and $H$, $S$ and $\tilde{S}$ depending on input $\bm{\theta}$ (e.g., interaction strengths). The $\bm{\theta}$ dependencies are implicitly assumed later to make notations concise. $\mathcal{A}$ could be response functions or scattering amplitudes Goldberger and Watson (1964); Efros (1985); Efros et al. (2007), [^1] For computing the response function of a ground state $|\Psi_{\text{gs}}\rangle$ induced by a multipole operator $O$ at real $E$, the compact sources are $O|\Psi_{\text{gs}}\rangle$. For scatterings, the sources could be compact as well Goldberger and Watson (1964); Efros (1985); Efros et al. (2007); also see later discussion of a two-body system. or an essential part of computing, e.g., optical potential Rotureau et al. (2017). They are basically $\langle\tilde{{S}}|\Psi\rangle$ or $\langle\tilde{\Psi}|{S}\rangle$ with $|\Psi\rangle$ and $\langle\tilde{\Psi}|$ satisfying the inhomogeneous Schrödinger equations:
$M|\Psi\rangle$ $=|{S}\rangle\,,\ \text{and}\ \ \langle\tilde{\Psi}|M=\langle\tilde{{S}}|\ .$ (3)
In the complex $E$ plane, $\mathcal{A}$ with a fixed $\bm{\theta}$ has poles at locations given by the bound-state eigenenergies. Branch points (i.e., thresholds) also exist at which kinetic phase spaces for the system’s fragmentations into subsystems start opening up; the associated branch cuts (BCTs) are typically along the real-$E$ axis due to $H$’s hermicity.
If $H$ is approximated by a finite Hermitian matrix, $\mathcal{A}$’s BCTs are discretized into poles located at the eigenenergies of the $H$-matrix—which are real. This gives a poor approximation of $\mathcal{A}$ because it contradicts the fact that $\mathcal{A}(E)$ must be continuous when varying $E$ along the real axis with $\mathrm{Im}\,E$ fixed to $0^{+}$ or $0^{-}$; $\mathcal{A}$ could change rapidly there, but that is generally due to the resonance poles on adjacent Riemann sheets, not the unphysical BCT poles.
The Type-I solution to this issue is to compute without discretizing the spectrum. It requires the continuum calculations (see, e.g., Glöckle (1983); Gloeckle et al. (1996); Deltuva et al. (2008, 2014); Lazauskas and Carbonell (2020); Marcucci et al. (2020); Navrátil et al. (2016); Navratil and Quaglioni (2022); Descouvemont and Baye (2010); Nielsen et al. (2001) for reviews) to handle wave functions’ oscillating asymptotics in coordinate space or singular functions in momentum space, when $E$ is real.
Type-II methods, or non-Hermitian quantum mechanics (NHQM) approaches Reinhardt (1982); Moiseyev (2011), move the BCTs below the real axis and discretize them. [^2] The BCT poles don’t necessarily line up along smooth BCT curves. Still, they must be far enough away from the real axis to separate them from the resonance poles. Unphysical poles are thus absent along the real axis. The resonance poles between the real axis and the new BCTs are now on the same Riemann sheet as the bound-state poles, meaning both types of poles can be seen in $\mathcal{A}$ simultaneously. Per Eq. (2), the resonances become $H$’s eigenstates. Methods of this type [^3] Some of these methods are in fact connected Afnan (1991). include integration contour deformation Glöckle (1983), complex scaling Reinhardt (1982); Moiseyev (2011); Myo et al. (2014); Lazauskas and Carbonell (2011); Lazauskas (2012); Papadimitriou and Vary (2015); Lazauskas (2015); Lazauskas and Carbonell (2020), and Berggren basis methods [^4] In the Berggren method, eigenenergies Michel et al. (2003); Michel and Płoszajczak (2021) were close to real energy axis, but they could be moved further away from the real axis by properly choosing single-particle momenta. Michel and Płoszajczak (2021). Here, the many-body basis states are typically the direct products of single-particle states.
Another challenge is exploring $\mathcal{A}$ in the space of $\bm{\theta}$ and $E$. It is usually infeasible to repeat the calculations many times. In fact, it may be unnecessary. According to model order reduction (MOR) studies Benner et al. (2017a, b, 2015), specifically those of the reduced basis methods (RBM) Hesthaven et al. (2015); Quarteroni et al. (2016); Duguet et al. (2023); Melendez et al. (2022); Drischler et al. (2022), the solution vector of an equation system typically moves in a low-dimensional subspace when varying the input parameters. By projecting the equation into the subspace, we create a reduced-order model or emulator, which can perform fast interpolation and extrapolation in the parameter space. The RBM emulators Duguet et al. (2023); Drischler et al. (2022); Melendez et al. (2022) are being studied for bound and resonance states Frame et al. (2018); Sarkar and Lee (2021, 2022); König et al. (2020); Demol et al. (2020); Ekström and Hagen (2019); Demol et al. (2021); Yoshida and Shimizu (2022); Anderson et al. (2022); Giuliani et al. (2023); Yapa et al. (2023) and general continuum states Furnstahl et al. (2020); Drischler et al. (2021); Melendez et al. (2021); Zhang and Furnstahl (2022); Bai and Ren (2021); Drischler and Zhang (2022); Melendez et al. (2022); Drischler et al. (2022); Bai (2022); Garcia et al. (2023); Odell et al. (2024). The latter emulators only work for the Type-I continuum calculations.
This work introduces a new NHQM approach that, via RBM-based subspace projections Melendez et al. (2022); Drischler et al. (2022); Duguet et al. (2023), moves BCTs away from the real energy axis and also emulate in $\bm{\theta}$. At the first step (or offline training stage), Eq. (3) is solved with a sample of $(E_{\alpha}^{tr},\bm{\theta}_{\alpha}^{tr})$ ($\alpha=1,,,N_{b}$, $\mathrm{Im}\,E_{\alpha}^{tr}>0$ and $\mathrm{Im}\,\bm{\theta}_{\alpha}^{tr}=0$). This step is the only computationally intensive component of the approach. The complex-$E$ solutions $|\Psi(E_{\alpha}^{tr},\bm{\theta}_{\alpha}^{tr})\rangle$ and $\langle\tilde{\Psi}(E_{\alpha}^{tr},\bm{\theta}_{\alpha}^{tr})|$, or simply $|\Psi_{\alpha}^{tr}\rangle$ and $\langle\tilde{\Psi}_{\alpha}^{tr}|$, satisfy bound-state-like boundary conditions Efros (1985); Efros et al. (2007); they also build in inter-particle correlations. At the next online emulation step, $|\Psi_{\alpha}^{tr}\rangle$ and $\langle\tilde{\Psi}_{\alpha}^{tr}|$ are used as a basis to form the subspace and generate general solutions,
$|\Psi\rangle$ $=c_{\alpha}|\Psi_{\alpha}^{tr}\rangle\,,\ \text{and}\ \ \langle\tilde{\Psi}|=\tilde{c}_{\alpha}\langle\tilde{\Psi}^{tr}_{\alpha}|\ .$ (4)
The convention of summing over repeated indices is used. To get $c_{\alpha}$ and $\tilde{c}_{\alpha}$ as functions of $E$ and $\bm{\theta}$, we plug the ansatz in Eq. (4) into a variational approach Pomraning (1965) for solving linear equations and get a small $N_{b}$-dimension linear system Zhang (2024): with $\begin{bmatrix}M\end{bmatrix}_{\alpha\beta}\equiv\langle\tilde{\Psi}^{tr}_{\alpha}|M|\Psi^{tr}_{\beta}\rangle$,
$\begin{bmatrix}M\end{bmatrix}_{\alpha\beta}c_{\beta}$ $=\langle\tilde{\Psi}^{tr}_{\alpha}|{S}\rangle\,,\ \text{and}\ \ \tilde{c}_{\beta}\begin{bmatrix}M\end{bmatrix}_{\beta\alpha}=\langle\tilde{{S}}|\Psi^{tr}_{\alpha}\rangle\ .$ (5)
The coefficients in these equations (e.g., $\begin{bmatrix}M\end{bmatrix}$) can be rapidly emulated when the parameter dependence in $M$ and ${S}$ and $\tilde{{S}}$ are factorized from the operators [^5] In this case (or with affine parameters), the operator matrix elements, which are expensive to compute, can be precalculated at the emulator training step and reused in the emulation step. The nonaffine dependences can be approximated by affine structures using various MOR methods (see, e.g., Benner et al. (2017a)). exactly (e.g., $E$ in $M$) or approximately. $\mathcal{A}$ can then be emulated for general $E$ and $\bm{\theta}$ values via $\begin{bmatrix}\mathcal{A}\end{bmatrix}=c_{\alpha}\langle\tilde{{S}}|\Psi^{tr}_{\alpha}\rangle=\tilde{c}_{\alpha}\langle\tilde{\Psi}^{tr}_{\alpha}|{S}\rangle$. I.e., we now have fast & accurate access to continuum physics (e.g., $\mathcal{A}$ at real $E$s) based on bound-state-like calculations.
In the following sections, we demonstrate the method in both two- and three-body systems—–importantly, the approach applies to general systems. Then, in the discussion section, we explain the method’s feasibility and its connections with existing methods, including using our method to emulate existing calculations based on the complex energy (CE) Schlessinger and Schwartz (1966); Schlessinger (1968a, b); McDonald and Nuttall (1969); Uzu et al. (2003); Deltuva and Fonseca (2012, 2013a, 2013b, 2014) and Lorentz integral transformation (LIT) methods Efros (1985); Efros et al. (1994, 2007); Orlandini et al. (2014); Sobczyk et al. (2021, 2024); Bonaiti et al. (2024). Afterward, future studies are discussed, followed by a brief summary.
Figure 1: The relative error (in its logarithm) of the emulated nonBorn term in two-body $s$-wave scattering $T$-matrix in the $E$’s complex plane. The training energies are evenly separated on the black solid line(s) and 10 MeV away from the real axis. See the text for more details.
Two-body demonstration.
A system of two particles mimicking nucleons in a $s$-wave channel is studied here. Each particle’s mass is $940$ MeV in natural units and the interaction $V$ is short-ranged. Let $|p_{in}\rangle$ be a plane wave state; both $|{S}\rangle$ and $|\tilde{{S}}\rangle$ are $V|p_{in}\rangle$ so that $\mathcal{A}(E)$ is the non-Born term [^6] The Born term costs much less to compute than the non-Born one. in the scattering $T$-matrix Efros (1985); Zhang (2024). Note we fix $\bm{\theta}$ (including $p_{in}$) and vary only $E$ here.
Figure 1 demonstrates the NHQM nature of the method, in particular, its dependence on $E_{\alpha}^{tr}$’s locations in the complex $E$ plane. It shows the emulation errors for $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ in the complex plane. In the top panel, $N_{b}=23$ training energies [^7] The $N_{b}$ dependence is studied in detail in Ref. Zhang (2024). were evenly spaced along the solid black line with $\mathrm{Im}\,E^{tr}_{\alpha}=10$ MeV. The error is the smallest ($\approx$ the training calculation errors) in the blue region, close to the training energies. When extrapolating, the error increases and diverges to infinity at the poles of $\begin{bmatrix}\mathcal{A}\end{bmatrix}$—see the dark orange dots. The curve formed by those pole dots can be considered as $\begin{bmatrix}\mathcal{A}\end{bmatrix}$’s discretized BCT originating from the $E=0$ branch point. Therefore, the BCT’s location—in this case, below the real axis—is directly connected to the emulation error pattern; the latter is controlled by the distribution of $E_{\alpha}^{tr}$.
This connection is also seen in the bottom panel, where ${E^{tr}_{\alpha}}^{\ast}$ are included in the training energies. There, the error pattern has a mirror symmetry with respect to the real axis, enforced by the symmetry of the training energy locations. As a result, the BCT is back on the real axis. The contrast between the two panels suggests that breaking the mirror symmetry of training energy locations with respect to the real axis forces the BCT away from the real axis, which, as discussed previously, is the basic feature of the NHQM methods.
To understand this point further, note that $\begin{bmatrix}\mathcal{A}\end{bmatrix}$’s poles are the poles of $c_{\alpha}$ and $\tilde{c}_{\alpha}$. Their locations are given by the solutions of the generalized eigenvalue problem with projected $H$ and norm matrices:
$\begin{bmatrix}H\end{bmatrix}_{\alpha\beta}=\langle\tilde{\Psi}_{\alpha}^{tr}|H|\Psi_{\beta}^{tr}\rangle\ \text{and}\ \begin{bmatrix}N\end{bmatrix}_{\alpha\beta}=\langle\tilde{\Psi}_{\alpha}^{tr}|\Psi_{\beta}^{tr}\rangle\ .$ (6)
Both are non-Hermitian generally, even with a Hermitian $H$ in the training calculations; they are complex symmetric if ${S}(\bm{\theta}_{\alpha})=\tilde{{S}}(\bm{\theta}_{\alpha})$ Zhang (2024). $\begin{bmatrix}H\end{bmatrix}$’s eigenenergies are thus complex, and $\begin{bmatrix}\mathcal{A}\end{bmatrix}$’s BCTs are off the real axis. But, as an exception, for the bottom panel of Fig. 1, the eigenvalue problem turns pseudo-Hermitian Mostafazadeh (2002a, b) with only real eigenenergies Zhang (2024).
Figure 2: Emulation in $E$ at fixed $\bm{\theta}$ for a three-body system with (top rows) and without (bottom rows) a bound dimer. The left panels show $\begin{bmatrix}\mathcal{A}\end{bmatrix}$’s eigenvalues ($\diamond$). Note different scales on the two sides of the x-axis there. The exact locations of bound or resonance states are also marked (see the legends). The insets zoom in on the regions around the branch points. The right panels compare the emulated and the exact calculations of $-\mathrm{Im}\,\mathcal{A}$ at real energies.
Three-body demonstrations.
We further illustrate our approach, including $\bm{\theta}$ emulation, for a three-body example. We employ the three-identical-boson model from Ref. Zhang and Furnstahl (2022) and apply our method to emulate the solutions of its Faddeev equations Glöckle (1983). Emulation of the original Schrödinger equation will be studied in the future. The equations can be cast into the form of Eq. (3), with
$|\Psi\rangle\equiv\begin{pmatrix}|\psi_{1}\rangle\\ |\psi_{4}\rangle\end{pmatrix}\ \mathrm{and}\ |\tilde{\Psi}\rangle\equiv\begin{pmatrix}|\tilde{\psi}_{1}\rangle\\ |\tilde{\psi}_{4}\rangle\end{pmatrix}\ ,$ (7)
and two-component $|{S}\rangle$ and $|\tilde{{S}}\rangle$. The sources are spatially localized and chosen so that $\langle\tilde{{S}}|\Psi\rangle$ and $\langle\tilde{\Psi}|{S}\rangle$ are the non-Born term in the particle-dimer scattering $T$-matrix Zhang (2024). With $H_{0}$ the kinetic energy operator, $V_{1}$ the pair-wise interactions with a coupling strength $\lambda$, $V_{4}$ the three-body interaction with a strength $\lambda_{4}$, and $\mathbb{P}$ the permutation operator Glöckle (1983), $M$ is
$\begin{pmatrix}E-H_{0}-V_{1}-V_{1}\mathbb{P}&-V_{1}\\ -3V_{4}&E-H_{0}-V_{4}\end{pmatrix}\equiv E-H\ .$ (8)
The full three-body wave function is $(1+\mathbb{P})|\psi_{1}\rangle+|\psi_{4}\rangle$. Note the model only considers $s$-wave dynamics Zhang and Furnstahl (2022); Zhang (2024).
We first extract the spectrum (i.e., eigenenergies) of $H(\bm{\theta})$ by emulating for $E$ while fixing $\bm{\theta}$ in $H$ and the sources. Figure 2 shows the spectra ($\diamond$ in left panels) and the related $-\mathrm{Im}\,\mathcal{A}$ (right panels) in two cases: the $H$ in the top row has a bound dimer with binding energy $B_{2}=10$ MeV, and the bottom has no bound dimers but a three-body resonance. The associated $\lambda_{4}$ and $\lambda$ (or $B_{2}$) are in the titles. Both emulators have the same sources and $N_{b}=48$ training energies with $\mathrm{Im}\,E_{\alpha}^{tr}=3$ MeV and $\mathrm{Re}\,E_{\alpha}^{tr}\in[-20,50]$ MeV.
[…]
Note that the BCT poles are exponentially clustered towards the branch point(s) in Figs. 2 (and 1). This distribution was also seen in recent works studying the so-called (near-)optimal rational approximation of a univariate function with branch points Nakatsukasa et al. (2018); Trefethen et al. (2021); Trefethen (2023); Nakatsukasa et al. (2023). This suggests a connection between the optimality of rational approximation and the effectiveness of the RBM subspace projection. See Ref. Zhang (2024) for detailed discussions.
[…]
The spectrum calculations can be emulated varying the $\bm{\theta}$ parameter in $H$, as shown in Fig. 3. The sources (fixed) and $\mathrm{Im}\,E_{\alpha}^{tr}$ are the same as in Fig. 2. In the top panels, $E$ and $\lambda_{4}$ are the emulation variables, while $\lambda$ is fixed. Two emulators with different $\lambda$ are trained (see the plot titles). We sample $N_{b}=60$ points in the $\mathrm{Re}\,E_{\alpha}^{tr}$-$\lambda_{4}$ space using Latin hypercube sampling (LHS) Tang (1993) with $\mathrm{Re}\,E_{\alpha}^{tr}\in[-20,50]$ MeV and $\lambda_{4}\in[-0.5,0.5]$. The spectrum emulated at a randomly chosen $\lambda_{4}$ is plotted for each emulator. In the bottom panels, $\lambda$ is included as an emulation variable. Thus, only one emulator is needed and trained, with $N_{b}=60$ training points sampled using LHS in the $\mathrm{Re}\,E_{\alpha}^{tr}$-$\lambda_{4}$-$\lambda$ space with $\lambda\in[-2.15,1]$ ($\mathrm{Re}\,E_{\alpha}^{tr}$ and $\lambda_{4}$ ranges unchanged). The emulator is tested at two parameter sets (see the plot titles). In the figure, the left panels have bound dimers and none for the right. These emulated spectra are qualitatively similar to those in Fig. 2, including the agreement between the emulated physical states and the exact results. These plotted spectra demonstrate the spectrum emulation capability of our method. However, deep into the complex plane (10s MeV below the real axis), the supposedly BCT eigenvalues are scattered somewhat. This is not problematic for emulating real-$E$ observables and near-axis resonances but can make isolating possible broad-resonance poles difficult. This phenomenon needs further studies.
Figure 4: The nonBorn term of the particle-dimer scattering $T$-matrix are emulated by two emulators with different $\mathrm{Im}\,E^{tr}_{\alpha}$. For each curve, $\lambda$ and thus $B_{2}$ are fixed (see the legend), but a sample of $\lambda_{4}$ values is chosen to test the emulators. The means of the sample of relative errors are plotted.
To emulate particle-dimer scattering amplitudes, we must vary the source parameters, including the interaction couplings and the scattering energy $E_{rel}$ between the particle and dimer, in addition to those in $H$. For on-shell scatterings, as computed in the emulation stage, the real-valued $E_{rel}$ is related to $E$ via $E_{rel}-B_{2}=E$. However, the training calculations treat them as separate emulation variables. Thus, we sample $N_{b}=128$ training points using LHS in the $\mathrm{Re}\,E$-$E_{rel}$-$\lambda$-$\lambda_{4}$ space, with $E_{rel}\in[0,60]$ MeV, $\lambda$ in a range in which $B_{2}\in[2,10]$ MeV, and $\mathrm{Re}\,E_{\alpha}^{tr}$ and $\lambda_{4}$ in the same ranges as for Fig. 3.
Two different emulators with $\mathrm{Im}\,E^{tr}_{\alpha}=3$ or $10$ are trained. They are checked against the exact on-shell $T$-matrices at a sample of 1000 points in the $E_{rel}$-$\lambda$-$\lambda_{4}$ space. Figure 4 demonstrates emulation performance by showing the relative errors of emulating the $T$-matrix’s nonBorn term versus $E_{rel}$. Each curve represents the mean of the relative errors from a sub-group of testing points with the same $B_{2}$ (see the legend) but different $\lambda_{4}$. The errors increase towards the particle-dimer threshold, similar to the behaviors of the rational approximations of functions near the branch points. Based on the comparisons, reducing $\mathrm{Im}\,E^{tr}_{\alpha}$ systemically improves the emulation, particularly around the branch points.
Discussion.
Our method has two components: the training calculations performed with complex $E_{\alpha}^{tr}$ and real $\bm{\theta}_{\alpha}^{tr}$, and the emulations that extrapolate the complex-$E$ training results to the other parts of the complex plane and interpolate (or extrapolate) in $\bm{\theta}$. The computing costs for emulation are low, but those of the training calculations increase with system size. Still, the complex-$E$ solutions have been computed in the continuum studies Carbonell et al. (2014) employing the CE Schlessinger and Schwartz (1966); Schlessinger (1968a, b); McDonald and Nuttall (1969); Uzu et al. (2003); Deltuva and Fonseca (2012, 2013a, 2013b, 2014) and LIT methods Efros (1985); Efros et al. (1994, 2007); Orlandini et al. (2014); Sobczyk et al. (2021, 2024); Bonaiti et al. (2024) for few- and many-body systems. Our method could thus apply to these systems as well.
To infer the real-$E$ results from the complex-$E$ solutions, the CE method uses rational-approximation based extrapolants Schlessinger and Schwartz (1966); Schlessinger (1968a, b); McDonald and Nuttall (1969); Uzu et al. (2003); Deltuva and Fonseca (2012, 2013a, 2013b, 2014) while LIT fits real-$E$ observables via integral transforms Efros (1985); Efros et al. (1994, 2007); Orlandini et al. (2014); Sobczyk et al. (2021, 2024); Bonaiti et al. (2024). In our method, the RBM emulator is used to realize this connection. It creates an NHQM approach that computes not only observables but also spectra in which bound and resonance states are separated from the BCT states. All three methods discussed here thus belong to the type-II methods mentioned earlier—an insight of this work.
Importantly, our method enables emulating the CE and LIT calculations. It provides fast interpolations of $\mathcal{A}$ at general $\bm{\theta}$ and $E$, given that $E$ is close to $E_{\alpha}^{tr}$. The complex-$E$ results at a given $\bm{\theta}$ can then be used as the inputs for the complex-$E$-to-real-$E$ procedures in CE and LIT. We thus create fast interpolations of the real-$E$ observables in the $\bm{\theta}$ space. See Ref. Zhang (2024) for details.
This work not only adds a new tool for studying continuum physics but also expands the MOR literature by applying the RBM method to project an operator with a continuum spectrum into a finite matrix; previous MOR literature Antoulas (2005) mainly focuses on dimension reduction of large matrices. Our study also suggests a connection between this RBM projection and (near-)optimal rational approximation of functions with branch points Nakatsukasa et al. (2018); Trefethen et al. (2021); Trefethen (2023); Nakatsukasa et al. (2023).
Further studies.
We derive essential understandings of the new method based on the training calculations with tiny errors (on the order of $10^{-12}$). Meanwhile, the numeric analytical continuation in $E$ is known to be sensitive to the errors in the data to be extrapolated Trefethen (2023). Such sensitivity will be addressed later. It will require the development of regularization methods, perhaps adapted to the specific many-body methods employed to perform the training calculations. For this purpose, the general behavior of the emulations exposed in this work, such as the distribution of eigenvalues, provides valuable prior knowledge for developing the regularization methods Hicks and Lee (2022).
Summary.
We apply the RBM method to emulate inhomogenous Schrödinger equations in a combined space of the complex $E$ plane and the parameters inside the Hamiltonian and the sources. It creates an NHQM method for extracting and emulating continuum states based on bound-state-like calculations. However, the method differs from previous NHQM methods in that it constructs the subspace by RBM-based projections, while the previous methods form a many-body basis with direct products of single-particle states. Good emulation performances are demonstrated in two and three-body systems as proof of principles. Moreover, the method can help emulate existing continuum calculations, such as those based on CE and LIT methods. Finally, the potential connection between the RBM’s effectiveness in mapping out the solution manifold and the optimality of the rational approximation should be further explored.
Acknowledgements.
Discussions with Dean Lee, Dick Furnstahl, Nobuo Hinohara, Chong Qi, Simin Wang, Alex Gnech, Bijaya Acharya, and Ante Ravlic are greatly appreciated. I also thank Dick for careful reading of the manuscript. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under the FRIB Theory Alliance Award No. DE-SC0013617 and under the STREAMLINE Collaboration Award No. DE-SC0024586.

### v2

@@
-
+Non-Hermitian Quantum Mechanics Approach for Extracting and Emulating Continuum Physics Based on Bound-State-Like Calculations
 Abstract
-This work develops a new method for computing a finite quantum system’s continuum states and observables by applying a subspace projection (or reduced basis) method used in model order reduction studies to “discretize” the system’s continuous spectrum.
-The method extracts the continuum physics from solving Schrödinger equations with bound-state-like boundary conditions and emulates this extraction in the space of the input parameters.
-This parameter emulation can readily be adapted to emulate other continuum calculations as well, e.g., those based on complex energy or Lorentz integral transform methods.
-Here, I give an overview of the key aspects of the formalism and some informative findings from numerical experimentation with two- and three-body systems, which indicates the non-Hermitian quantum mechanics nature of the method.
-A potential connection with (near-)optimal rational approximation studied in Math literature is also discussed.
-Further details are provided in a separate paper.
+This work introduces a unified emulation framework for studying continuum physics in finite quantum systems.
+Using a reduced basis method, we construct powerful emulators for the inhomogeneous Schrödinger equation that operate in a combined parameter space of complex energy ($E$) and other inputs ($\bm{\theta}$).
+Within the space, the emulators simultaneously perform analytical continuation in EE—extracting continuum physics from numerically simpler bound-state-like calculations—and interpolate this entire process across $\bm{\theta}$.
+This yields a small, non-Hermitian system whose properties (e.g., resonances and scattering observables) can be rapidly predicted for any $\bm{\theta}$.
+Crucially, the complex-$E$ emulation provides a pathway to compute continuum observables for complex systems where advanced bound-state methods exist but direct continuum calculations are yet to be developed, while the $\bm{\theta}$-emulation enables rapid parameter-space exploration and can be adapted to accelerate other existing continuum calculations.
+Demonstrations with two- and three-body systems highlight the method’s effectiveness and suggest its connection to (near-)optimal rational approximation.
+This Letter presents the key results, with further details reserved for a companion paper.
 Introduction.
@@
 $\mathcal{A}(E,\bm{\theta})\equiv\left\langle\tilde{S}(\bm{\theta})\left|\left[M(E,\bm{\theta})\right]^{-1}\right|S(\bm{\theta})\right\rangle\ ,$ (1)
-with $M(E,\bm{\theta})\equiv E-H(\bm{\theta})$, and $H$, $S$ and $\tilde{S}$ depending on input $\bm{\theta}$ (e.g., interaction strengths).
-The $\bm{\theta}$ dependencies are implicitly assumed later to make notations concise.
-$\mathcal{A}$ could be response functions or scattering amplitudes Goldberger and Watson (1964); Efros (1985); Efros et al.
-(2007), [^1] For computing the response function of a ground state $|\Psi_{\text{gs}}\rangle$ induced by a multipole operator $O$ at real $E$, the compact sources are $O|\Psi_{\text{gs}}\rangle$.
-For scatterings, the sources could be compact as well Goldberger and Watson (1964); Efros (1985); Efros et al.
-(2007); also see later discussion of a two-body system.
-or an essential part of computing, e.g., optical potential Rotureau et al.
-(2017).
-They are basically $\langle\tilde{{S}}|\Psi\rangle$ or $\langle\tilde{\Psi}|{S}\rangle$ with $|\Psi\rangle$ and $\langle\tilde{\Psi}|$ satisfying the inhomogeneous Schrödinger equations:
-$M|\Psi\rangle$ $=|{S}\rangle\,,\ \text{and}\ \ \langle\tilde{\Psi}|M=\langle\tilde{{S}}|\ .$ (3)
+[…]
+[…]
 In the complex $E$ plane, $\mathcal{A}$ with a fixed $\bm{\theta}$ has poles at locations given by the bound-state eigenenergies.
-Branch points (i.e., thresholds) also exist at which kinetic phase spaces for the system’s fragmentations into subsystems start opening up; the associated branch cuts (BCTs) are typically along the real-$E$ axis due to $H$’s hermicity.
+Branch points (i.e., thresholds) also exist at which kinetic phase spaces for the system’s fragmentation into subsystems start opening up; the associated branch cuts (BCTs) are typically along the real-$E$ axis due to $H$’s Hermiticity [4].
 If $H$ is approximated by a finite Hermitian matrix, $\mathcal{A}$’s BCTs are discretized into poles located at the eigenenergies of the $H$-matrix—which are real.
@@
 The type-I solution to this issue is to compute without discretizing the spectrum.
-It requires the continuum calculations (see, e.g., Glöckle (1983); Gloeckle et al.
-(1996); Deltuva et al.
-(2008, 2014); Lazauskas and Carbonell (2020); Marcucci et al.
-(2020); Navrátil et al.
-(2016); Navratil and Quaglioni (2022); Descouvemont and Baye (2010); Nielsen et al.
-(2001) for reviews) to handle wave functions’ oscillating asymptotics in coordinate space or singular functions in momentum space, when $E$ is real.
-Type-II methods, or non-Hermitian quantum mechanics (NHQM) approaches Reinhardt (1982); Moiseyev (2011), move the BCTs below the real axis and discretize them.
-[^2] The BCT poles don’t necessarily line up along smooth BCT curves.
-Still, they must be far enough away from the real axis to separate them from the resonance poles.
-Unphysical poles are thus absent along the real axis.
-The resonance poles between the real axis and the new BCTs are now on the same Riemann sheet as the bound-state poles, meaning both types of poles can be seen in $\mathcal{A}$ simultaneously.
+It requires the calculations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15] to handle wave functions’ complicated oscillating asymptotics in coordinate space or singular functions in momentum space, when $E$ is real.
+The type-II methods, or non-Hermitian quantum mechanics (NHQM) approaches [16, 17, 18], move the BCTs below the real axis and discretize them [16, 17]   [^2] The BCT poles don’t necessarily line up along smooth BCT curves.
+Still, they must be far enough away from the real axis to separate them from the resonance poles..
+Unphysical poles are absent on the real axis.
+The resonance poles above the new BCTs are now on the same Riemann sheet as the bound-state’s, meaning both can be seen in $\mathcal{A}$ simultaneously.
 Per Eq.
 (1), the resonances become $H$’s eigenstates.
-Methods of this type [^3] Some of these methods are in fact connected Afnan (1991).
-include integration contour deformation Glöckle (1983), complex scaling Reinhardt (1982); Moiseyev (2011); Myo et al.
-(2014); Lazauskas and Carbonell (2011); Lazauskas (2012); Papadimitriou and Vary (2015); Lazauskas (2015); Lazauskas and Carbonell (2020), and Berggren basis methods [^4] In the Berggren method, eigenenergies Michel et al.
-(2003); Michel and Płoszajczak (2021) were close to real energy axis, but they could be moved further away from the real axis by properly choosing single-particle momenta.
-Michel and Płoszajczak (2021).
-Here, the many-body basis states are typically the direct products of single-particle states.
+Methods of this type   [^3] Some of these methods are in fact connected [19].
+include integration contour deformation [6], complex scaling [17, 18, 20, 21, 22, 23, 24, 10], and Berggren basis methods [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
 Another challenge is exploring $\mathcal{A}$ in the space of $\bm{\theta}$ and $E$.
@@
 In fact, it may be unnecessary.
-According to model order reduction (MOR) studies Benner et al.
-(2017a, b, 2015), specifically those of the reduced basis methods (RBM) Hesthaven et al.
-(2015); Quarteroni et al.
-(2016); Duguet et al.
-(2023); Melendez et al.
-(2022); Drischler et al.
-(2022), the solution vector of an equation system typically moves in a low-dimensional subspace when varying the input parameters.
-By projecting the equation into the subspace, we create a reduced-order model or emulator, which can perform fast interpolation and extrapolation in the parameter space.
-The RBM emulators Duguet et al.
-(2023); Drischler et al.
-(2022); Melendez et al.
-(2022) are being studied for bound and resonance states Frame et al.
-(2018); Sarkar and Lee (2021, 2022); König et al.
-(2020); Demol et al.
-(2020); Ekström and Hagen (2019); Demol et al.
-(2021); Yoshida and Shimizu (2022); Anderson et al.
-(2022); Giuliani et al.
-(2023); Yapa et al.
-(2023) and general continuum states Furnstahl et al.
-(2020); Drischler et al.
-(2021); Melendez et al.
-(2021); Zhang and Furnstahl (2022); Bai and Ren (2021); Drischler and Zhang (2022); Melendez et al.
-(2022); Drischler et al.
-(2022); Bai (2022); Garcia et al.
-(2023); Odell et al.
-(2024).
-The latter emulators only work for the Type-I continuum calculations.
-This work introduces a new NHQM approach that, via RBM-based subspace projections Melendez et al.
-(2022); Drischler et al.
-(2022); Duguet et al.
-(2023), moves BCTs away from the real energy axis and also emulate in $\bm{\theta}$.
-At the first step (or offline training stage), Eq.
-(3) is solved with a sample of $(E_{\alpha}^{tr},\bm{\theta}_{\alpha}^{tr})$ ($\alpha=1,,,N_{b}$, $\mathrm{Im}\,E_{\alpha}^{tr}>0$ and $\mathrm{Im}\,\bm{\theta}_{\alpha}^{tr}=0$).
+According to the reduced basis methods (RBM) [36, 37, 38, 39, 40] in model order reduction (MOR) [41, 42, 43], the solution vector of an equation system, when varying a number ($N_{p}$) of its inputs, typically moves in a subspace with a low dimensionality (scaling mildly with $N_{p}$).
+Projecting the equations onto that subspace creates a smaller equation system, known as reduced-order model—a particular type of emulator or surrogate model which accurately and rapidly explores the parameter space [39, 40, 38].
+Such emulators are being studied for bound and resonance states [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] and continuum states  [^4] Also see Ref.
+[55] for perturbation-based scattering emulators.[56, 57, 58, 59, 60, 61, 39, 40, 62, 63, 64, 65, 66].
+This work introduces a new NHQM approach that, via RBM-based subspace projections [39, 40, 38], simultaneously (1) moves $\mathcal{A}(E,\bm{\theta})$’s BCTs below the real-$E$ axis in the complex-$E$ plane and analytically continues $\mathcal{A}$ from above real-$E$ axis downward to the real-$E$ axis and further into the resonance region and (2) emulates the analytical continuation process in the space of real $\bm{\theta}$.
+Specifically, at the first offline training stage, Eq.
+(2) is solved at a sample of $(E_{\alpha}^{\mathrm{tr}},\bm{\theta}_{\alpha}^{\mathrm{tr}})$ ($\alpha=1,,,N_{b}$) with $\mathrm{Im}\,E_{\alpha}^{\mathrm{tr}}>0$.
 This step is the only computationally intensive component of the approach.
-The complex-$E$ solutions $|\Psi(E_{\alpha}^{tr},\bm{\theta}_{\alpha}^{tr})\rangle$ and $\langle\tilde{\Psi}(E_{\alpha}^{tr},\bm{\theta}_{\alpha}^{tr})|$, or simply $|\Psi_{\alpha}^{tr}\rangle$ and $\langle\tilde{\Psi}_{\alpha}^{tr}|$, satisfy bound-state-like boundary conditions Efros (1985); Efros et al.
-(2007); they also build in inter-particle correlations.
-At the next online emulation step, $|\Psi_{\alpha}^{tr}\rangle$ and $\langle\tilde{\Psi}_{\alpha}^{tr}|$ are used as a basis to form the subspace and generate general solutions,
-$|\Psi\rangle$ $=c_{\alpha}|\Psi_{\alpha}^{tr}\rangle\,,\ \text{and}\ \ \langle\tilde{\Psi}|=\tilde{c}_{\alpha}\langle\tilde{\Psi}^{tr}_{\alpha}|\ .$ (4)
+The solutions $|\Psi(E_{\alpha}^{\mathrm{tr}},\bm{\theta}_{\alpha}^{\mathrm{tr}})\rangle$ and $\langle\tilde{\Psi}(E_{\alpha}^{\mathrm{tr}},\bm{\theta}_{\alpha}^{\mathrm{tr}})|$, or simply $|\Psi_{\alpha}^{\mathrm{tr}}\rangle$ and $\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|$, satisfy bound-state-like boundary conditions [2, 3]—which are much simpler to deal with numerically than the oscillating asymptotics in the real-$E$ cases.
+At the next online emulation step, $|\Psi_{\alpha}^{\mathrm{tr}}\rangle$ and $\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|$ are used as a subspace basis to construct general solutions:
+$|\Psi\rangle$ $=c_{\alpha}|\Psi_{\alpha}^{\mathrm{tr}}\rangle\,\ \text{and}\ \ \langle\tilde{\Psi}|=\tilde{c}_{\alpha}\langle\tilde{\Psi}^{\mathrm{tr}}_{\alpha}|\ .$ (3)
 The convention of summing over repeated indices is used.
+In this work, for simplicity, we typically over-sample the training points, though a singular value decomposition (SVD) of the norm matrix [defined in Eq.
+(5)] reveals a much smaller effective rank ($N_{\mathrm{eff}}\ll N_{b}$).
+This redundancy points to future optimizations with greedy algorithms that pick training points in a more strategic and efficient way  [46, 66, 37].
 To get $c_{\alpha}$ and $\tilde{c}_{\alpha}$ as functions of $E$ and $\bm{\theta}$, we plug the ansatz in Eq.
-(4) into a variational approach Pomraning (1965) for solving linear equations and get a small $N_{b}$-dimension linear system Zhang (2024): with $\begin{bmatrix}M\end{bmatrix}_{\alpha\beta}\equiv\langle\tilde{\Psi}^{tr}_{\alpha}|M|\Psi^{tr}_{\beta}\rangle$,
-$\begin{bmatrix}M\end{bmatrix}_{\alpha\beta}c_{\beta}$ $=\langle\tilde{\Psi}^{tr}_{\alpha}|{S}\rangle\,,\ \text{and}\ \ \tilde{c}_{\beta}\begin{bmatrix}M\end{bmatrix}_{\beta\alpha}=\langle\tilde{{S}}|\Psi^{tr}_{\alpha}\rangle\ .$ (5)
-The coefficients in these equations (e.g., $\begin{bmatrix}M\end{bmatrix}$) can be rapidly emulated when the parameter dependence in $M$ and ${S}$ and $\tilde{{S}}$ are factorized from the operators [^5] In this case (or with affine parameters), the operator matrix elements, which are expensive to compute, can be precalculated at the emulator training step and reused in the emulation step.
-The nonaffine dependences can be approximated by affine structures using various MOR methods (see, e.g., Benner et al.
-(2017a)).
-exactly (e.g., $E$ in $M$) or approximately.
-$\mathcal{A}$ can then be emulated for general $E$ and $\bm{\theta}$ values via $\begin{bmatrix}\mathcal{A}\end{bmatrix}=c_{\alpha}\langle\tilde{{S}}|\Psi^{tr}_{\alpha}\rangle=\tilde{c}_{\alpha}\langle\tilde{\Psi}^{tr}_{\alpha}|{S}\rangle$.
-I.e., we now have fast & accurate access to continuum physics (e.g., $\mathcal{A}$ at real $E$s) based on bound-state-like calculations.
+(3) into a variational approach [67] for solving linear equations and get a small $N_{b}$-dimension linear system [4]: with $\begin{bmatrix}M\end{bmatrix}_{\alpha\beta}\equiv\langle\tilde{\Psi}^{\mathrm{tr}}_{\alpha}|M|\Psi^{\mathrm{tr}}_{\beta}\rangle$,
+$\begin{bmatrix}M\end{bmatrix}_{\alpha\beta}c_{\beta}$ $=\langle\tilde{\Psi}^{\mathrm{tr}}_{\alpha}|{S}\rangle\,\ \text{and}\ \ \tilde{c}_{\beta}\begin{bmatrix}M\end{bmatrix}_{\beta\alpha}=\langle\tilde{{S}}|\Psi^{\mathrm{tr}}_{\alpha}\rangle\ .$ (4)
+The coefficients in these equations (e.g., $\begin{bmatrix}M\end{bmatrix}$) can be rapidly emulated when the parameter dependence in $M$ and S{S} and $\tilde{{S}}$ are factorized from the operators exactly (e.g., $E$ in $M$) or approximately  [^5] Detailed discussion on this point can be found in e.g., Refs.
+[40, 64], concerning affine vs non-affine parameters..
+$\mathcal{A}$ can then be emulated for general $E$ and $\bm{\theta}$ values via $\begin{bmatrix}\mathcal{A}\end{bmatrix}=c_{\alpha}\langle\tilde{{S}}|\Psi^{\mathrm{tr}}_{\alpha}\rangle=\tilde{c}_{\alpha}\langle\tilde{\Psi}^{\mathrm{tr}}_{\alpha}|{S}\rangle$.
+Note that the approach’s reliance on bound-state-like calculations provides a key strategic advantage [18, 3, 68, 69]: leveraging advanced bound-state methods—which are generally applicable to larger systems than their continuum counterparts—to expand the reach of continuum calculations.
+In short, we now have fast & accurate access to continuum physics, e.g., $\mathcal{A}(E,\theta)$, at different real values of $E$ and $\theta$, based on bound-state-like calculations.
 In the following sections, we demonstrate the method in both two- and three-body systems—–importantly, the approach applies to general systems.
-Then, in the discussion section, we explain the method’s feasibility and its connections with existing methods, including using our method to emulate existing calculations based on the complex energy (CE) Schlessinger and Schwartz (1966); Schlessinger (1968a, b); McDonald and Nuttall (1969); Uzu et al.
-(2003); Deltuva and Fonseca (2012, 2013a, 2013b, 2014) and Lorentz integral transformation (LIT) methods Efros (1985); Efros et al.
-(1994, 2007); Orlandini et al.
-(2014); Sobczyk et al.
-(2021, 2024); Bonaiti et al.
-(2024).
+Multiple ways of varying $E$ and or $\bm{\theta}$ will be used to realize the method’s different functionalities.
+Then, the discussion section offers a broad perspective about the method, including its connections with the existing complex energy (CE) [70, 71, 72, 73, 74, 75, 76, 77, 78] and Lorentz integral transformation (LIT) [2, 79, 3, 69, 80, 81, 82] methods.
 Afterward, future studies are discussed, followed by a brief summary.
-Figure 1: The relative error (in its logarithm) of the emulated nonBorn term in two-body $s$-wave scattering $T$-matrix in the $E$’s complex plane.
-The training energies are evenly separated on the black solid line(s) and 10 MeV away from the real axis.
-See the text for more details.
+A detailed description of the work is provided in Ref.
+[4].
+Source codes for generating the results of this work can be accessed via the companion website [83].
+Figure 1: The relative error (in its logarithm) of an emulated resolvent matrix element for a two-body system in the complex-$E$ plane.
+The training energies are evenly separated on the black solid line(s) which are 10 MeV away from the real axis.
 Two-body demonstration.
-A system of two particles mimicking nucleons in a $s$-wave channel is studied here.
-Each particle’s mass is $940$ MeV in natural units and the interaction $V$ is short-ranged.
-Let $|p_{in}\rangle$ be a plane wave state; both $|{S}\rangle$ and $|\tilde{{S}}\rangle$ are $V|p_{in}\rangle$ so that $\mathcal{A}(E)$ is the non-Born term [^6] The Born term costs much less to compute than the non-Born one.
-in the scattering $T$-matrix Efros (1985); Zhang (2024).
-Note we fix $\bm{\theta}$ (including $p_{in}$) and vary only $E$ here.
-Figure 1 demonstrates the NHQM nature of the method, in particular, its dependence on $E_{\alpha}^{tr}$’s locations in the complex $E$ plane.
+We study a two-nucleon-like system in their $s$-wave channel, with a short-ranged interaction $V$ (see Ref.
+[4] for more details).
+Particle mass is $940$ MeV in natural units.
+Let $|p_{in}\rangle$ be a plane wave state, and $|{S}\rangle$ and $|\tilde{{S}}\rangle$ as $V|p_{in}\rangle$.
+We fix $H$ and sources (including $p_{in}$) and vary $E$.
+With such sources, $\mathcal{A}$ becomes the non-Born term in the scattering $T$-matrix [2, 4], but $\mathrm{Im}\,\mathcal{A}$ also mimics response functions.
+Note that all the results of this work are based on training calculations with tiny errors ($\approx 10^{-12}$), allowing us to derive essential understandings of the approach.
+Figure 1 demonstrates the NHQM nature of the method, in particular, its dependence on $E_{\alpha}^{\mathrm{tr}}$’s locations in the complex-$E$ plane.
 It shows the emulation errors for $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ in the complex plane.
-In the top panel, $N_{b}=23$ training energies [^7] The $N_{b}$ dependence is studied in detail in Ref.
-Zhang (2024).
-were evenly spaced along the solid black line with $\mathrm{Im}\,E^{tr}_{\alpha}=10$ MeV.
+In the top panel, $N_{b}=23$ training energies were evenly spaced along the solid black line with $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}=10$ MeV (note that the effective basis size $N_{\mathrm{eff}}\approx 10$, see Ref.
+[4] for more details).
 The error is the smallest ($\approx$ the training calculation errors) in the blue region, close to the training energies.
@@
 The curve formed by those pole dots can be considered as $\begin{bmatrix}\mathcal{A}\end{bmatrix}$’s discretized BCT originating from the $E=0$ branch point.
-Therefore, the BCT’s location—in this case, below the real axis—is directly connected to the emulation error pattern; the latter is controlled by the distribution of $E_{\alpha}^{tr}$.
-This connection is also seen in the bottom panel, where ${E^{tr}_{\alpha}}^{\ast}$ are included in the training energies.
+Therefore, the BCT’s location—in this case, below the real axis—is directly connected to the emulation error pattern; the latter is controlled by the locations of $E_{\alpha}^{\mathrm{tr}}$.
+This connection is also seen in the bottom panel, where $E^{\mathrm{tr}}_{\alpha}$$^{\ast}$ are included in the training energies.
 There, the error pattern has a mirror symmetry with respect to the real axis, enforced by the symmetry of the training energy locations.
@@
 Their locations are given by the solutions of the generalized eigenvalue problem with projected $H$ and norm matrices:
-$\begin{bmatrix}H\end{bmatrix}_{\alpha\beta}=\langle\tilde{\Psi}_{\alpha}^{tr}|H|\Psi_{\beta}^{tr}\rangle\ \text{and}\ \begin{bmatrix}N\end{bmatrix}_{\alpha\beta}=\langle\tilde{\Psi}_{\alpha}^{tr}|\Psi_{\beta}^{tr}\rangle\ .$ (6)
-Both are non-Hermitian generally, even with a Hermitian $H$ in the training calculations; they are complex symmetric if ${S}(\bm{\theta}_{\alpha})=\tilde{{S}}(\bm{\theta}_{\alpha})$ Zhang (2024).
+$\begin{bmatrix}H\end{bmatrix}_{\alpha\beta}=\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|H|\Psi_{\beta}^{\mathrm{tr}}\rangle\ \text{and}\ \begin{bmatrix}N\end{bmatrix}_{\alpha\beta}=\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|\Psi_{\beta}^{\mathrm{tr}}\rangle\ .$ (5)
+Both are non-Hermitian generally, even with a Hermitian $H$ in the training calculations; they are complex symmetric if ${S}(\bm{\theta}_{\alpha})$ and $\tilde{{S}}(\bm{\theta}_{\alpha})$ are equal and invariant under time-reversal transformation.
 $\begin{bmatrix}H\end{bmatrix}$’s eigenenergies are thus complex, and $\begin{bmatrix}\mathcal{A}\end{bmatrix}$’s BCTs are off the real axis.
 But, as an exception, for the bottom panel of Fig.
-1, the eigenvalue problem turns pseudo-Hermitian Mostafazadeh (2002a, b) with only real eigenenergies Zhang (2024).
+1, the eigenvalue problem turns Hermitian with only real eigenenergies.
+See Ref.
+[4] for detailed discussions.
 Figure 2: Emulation in $E$ at fixed $\bm{\theta}$ for a three-body system with (top rows) and without (bottom rows) a bound dimer.
 The left panels show $\begin{bmatrix}\mathcal{A}\end{bmatrix}$’s eigenvalues ($\diamond$).
-Note different scales on the two sides of the x-axis there.
+Note the different scales on the two sides of the x-axis.
 The exact locations of bound or resonance states are also marked (see the legends).
 The insets zoom in on the regions around the branch points.
-The right panels compare the emulated and the exact calculations of $-\mathrm{Im}\,\mathcal{A}$ at real energies.
+The right panels compare the emulated and the exact calculations of $\mathrm{Im}\,\mathcal{A}$ at real energies.
 Three-body demonstrations.
-We further illustrate our approach, including $\bm{\theta}$ emulation, for a three-body example.
-We employ the three-identical-boson model from Ref.
-Zhang and Furnstahl (2022) and apply our method to emulate the solutions of its Faddeev equations Glöckle (1983).
-Emulation of the original Schrödinger equation will be studied in the future.
+We illustrate our method, including its $\bm{\theta}$-emulation, for a $s$-wave interacting three-boson system studied previously in Ref.
+[59], by emulating the solutions of Faddeev equations [6].
+Emulation in terms of the Schrödinger equation will be studied in the future.
 The equations can be cast into the form of Eq.
 (2), with
-$|\Psi\rangle\equiv\begin{pmatrix}|\psi_{1}\rangle\\ |\psi_{4}\rangle\end{pmatrix}\ \mathrm{and}\ |\tilde{\Psi}\rangle\equiv\begin{pmatrix}|\tilde{\psi}_{1}\rangle\\ |\tilde{\psi}_{4}\rangle\end{pmatrix}\ ,$ (7)
+$|\Psi\rangle\equiv\begin{pmatrix}|\psi_{1}\rangle\\ |\psi_{4}\rangle\end{pmatrix}\ \mathrm{and}\ |\tilde{\Psi}\rangle\equiv\begin{pmatrix}|\tilde{\psi}_{1}\rangle\\ |\tilde{\psi}_{4}\rangle\end{pmatrix}\$, (6)
 and two-component $|{S}\rangle$ and $|\tilde{{S}}\rangle$.
-The sources are spatially localized and chosen so that $\langle\tilde{{S}}|\Psi\rangle$ and $\langle\tilde{\Psi}|{S}\rangle$ are the non-Born term in the particle-dimer scattering $T$-matrix Zhang (2024).
-With $H_{0}$ the kinetic energy operator, $V_{1}$ the pair-wise interactions with a coupling strength $\lambda$, $V_{4}$ the three-body interaction with a strength $\lambda_{4}$, and $\mathbb{P}$ the permutation operator Glöckle (1983), $M$ is
+The sources are spatially localized and chosen so that $\langle\tilde{{S}}|\Psi\rangle$ and $\langle\tilde{\Psi}|{S}\rangle$ are the non-Born term in the particle-dimer scattering $T$-matrix [4].
+With $H_{0}$ the kinetic energy operator, $V_{1}$ the pair-wise interactions with a coupling strength $\lambda$, $V_{4}$ the three-body interaction with a strength $\lambda_{4}$, and $\mathbb{P}$ the permutation operator [6], $M$ is
 $\begin{pmatrix}E-H_{0}-V_{1}-V_{1}\mathbb{P}&-V_{1}\\ -3V_{4}&E-H_{0}-V_{4}\end{pmatrix}\equiv E-H\ .$ (7)
-The full three-body wave function is $(1+\mathbb{P})|\psi_{1}\rangle+|\psi_{4}\rangle$. Note the model only considers $s$-wave dynamics Zhang and Furnstahl (2022); Zhang (2024).
-We first extract the spectrum (i.e., eigenenergies) of $H(\bm{\theta})$ by emulating for $E$ while fixing $\bm{\theta}$ in $H$ and the sources.
-Figure 2 shows the spectra ($\diamond$ in left panels) and the related $-\mathrm{Im}\,\mathcal{A}$ (right panels) in two cases: the $H$ in the top row has a bound dimer with binding energy $B_{2}=10$ MeV, and the bottom has no bound dimers but a three-body resonance.
+Below, we (1) vary $E$ alone to compute spectrum (i.e., eigenenergies) and $\mathcal{A}$ for a fixed $H$ and sources (as done in the two-body demonstration); (2) alter $E$ and the $\bm{\theta}$ inside $H(\bm{\theta})$ to emulate the spectra and $\mathcal{A}$ for different Hamiltonians; and (3) vary $E$ and the $\bm{\theta}$ in both $H$ and sources to gain the full set of functionalities, e.g., computing scattering amplitudes and emulating them in $\bm{\theta}$.
+We first extract the spectrum of $H(\bm{\theta})$ by emulating in $E$ while fixing $\bm{\theta}$ in $H$ and the sources.
+Figure 2 shows the spectra ($\diamond$ in left panels) and the related $\mathrm{Im}\,\mathcal{A}$ (right panels) in two cases: the $H$ in the top row has a bound dimer with binding energy $B_{2}=10$ MeV, and the bottom has no bound dimers but a three-body resonance.
 The associated $\lambda_{4}$ and $\lambda$ (or $B_{2}$) are in the titles.
-Both emulators have the same sources and $N_{b}=48$ training energies with $\mathrm{Im}\,E_{\alpha}^{tr}=3$ MeV and $\mathrm{Re}\,E_{\alpha}^{tr}\in[-20,50]$ MeV.
+Both emulators have the same sources and $N_{b}=48$ training energies with $\mathrm{Im}\,E_{\alpha}^{\mathrm{tr}}=3$ MeV and $\mathrm{Re}\,E_{\alpha}^{\mathrm{tr}}\in[-20,50]$ MeV ($N_{\mathrm{eff}}\approx 15$ for both cases).
+The left panels in Fig.
+2 again demonstrate the NHQM nature of the method: most eigenenergies form off-axis curves, representing $\begin{bmatrix}\mathcal{A}\end{bmatrix}$’s discretized BCTs which start from the branch point(s) on the real axis (exact values marked by black vertical lines).
+In the top left panel, there are two branch points corresponding to particle-dimer and three-particle thresholds, while only a single branch point exists in the bottom left.
 […]
-Note that the BCT poles are exponentially clustered towards the branch point(s) in Figs.
-2 (and 1).
-This distribution was also seen in recent works studying the so-called (near-)optimal rational approximation of a univariate function with branch points Nakatsukasa et al.
-(2018); Trefethen et al.
-(2021); Trefethen (2023); Nakatsukasa et al.
-(2023).
-This suggests a connection between the optimality of rational approximation and the effectiveness of the RBM subspace projection.
-See Ref.
-Zhang (2024) for detailed discussions.
+Importantly, the pattern in which the BCT states and isolated bound or resonance states are distributed in Fig.
+2 helps us identify the physical states [4], considering that we can’t order the complex eigenenergies to identify states as done in dealing with real spectra.
+Moreover, the BCT poles are exponentially clustered towards the branch point(s) in Figs.
+2 (and 1), mirroring that found in recent studies of (near-)optimal rational approximation (ORA) for functions with branch points [84, 85, 86, 87].
+Such similarities [4] suggest a connection between the two methods.
+The robustness of ORA [84, 87] thus lends support to our complex-$E$ emulation as an advanced analytical continuation tool.
+Finally, in the right panels of Fig.
+2, the emulated and exact calculations of $\mathrm{Im}\,\mathcal{A}$ at real energies are very close, including near thresholds.
+Note that $\mathrm{Im}\,\mathcal{A}$ mimics typical response functions.
 […]
-The spectrum calculations can be emulated varying the $\bm{\theta}$ parameter in $H$, as shown in Fig.
-3.
-The sources (fixed) and $\mathrm{Im}\,E_{\alpha}^{tr}$ are the same as in Fig.
-2.
-In the top panels, $E$ and $\lambda_{4}$ are the emulation variables, while $\lambda$ is fixed.
-Two emulators with different $\lambda$ are trained (see the plot titles).
-We sample $N_{b}=60$ points in the $\mathrm{Re}\,E_{\alpha}^{tr}$-$\lambda_{4}$ space using Latin hypercube sampling (LHS) Tang (1993) with $\mathrm{Re}\,E_{\alpha}^{tr}\in[-20,50]$ MeV and $\lambda_{4}\in[-0.5,0.5]$.
-The spectrum emulated at a randomly chosen $\lambda_{4}$ is plotted for each emulator.
-In the bottom panels, $\lambda$ is included as an emulation variable.
-Thus, only one emulator is needed and trained, with $N_{b}=60$ training points sampled using LHS in the $\mathrm{Re}\,E_{\alpha}^{tr}$-$\lambda_{4}$-$\lambda$ space with $\lambda\in[-2.15,1]$ ($\mathrm{Re}\,E_{\alpha}^{tr}$ and $\lambda_{4}$ ranges unchanged).
-The emulator is tested at two parameter sets (see the plot titles).
-In the figure, the left panels have bound dimers and none for the right.
+Now, can we emulate the spectrum calculations for different $H$?
+Yes, if we vary $E$ and the $\bm{\theta}$ inside $H(\bm{\theta})$.
+Here, a single emulator is trained to emulate in $E$, $\lambda_{4}$ and $\lambda$.
+We fix the sources and $\mathrm{Im}\,E_{\alpha}^{\mathrm{tr}}$ to be the same as in Fig.
+2, and then sample $N_{b}=60$ training points ($N_{\mathrm{eff}}\approx 43$) in the $(\mathrm{Re}\,E_{\alpha}^{\mathrm{tr}},\lambda_{4},\lambda)$ space using Latin hypercube sampling (LHS) [88] with $\mathrm{Re}\,E_{\alpha}^{\mathrm{tr}}\in[-20,50]$ MeV, $\lambda_{4}\in[-0.5,0.5]$, and $\lambda\in[-2.15,1]$.
+The emulator is then tested at two instances, as shown in Fig.
+3, corresponding to two distinctive systems: one has a bound dimer in the left, while none in the right (see the panel titles for the parameter values).
 These emulated spectra are qualitatively similar to those in Fig.
 2, including the agreement between the emulated physical states and the exact results.
-These plotted spectra demonstrate the spectrum emulation capability of our method.
+The results thus demonstrate the spectrum emulation capability of our method.
 However, deep into the complex plane (10s MeV below the real axis), the supposedly BCT eigenvalues are scattered somewhat.
-This is not problematic for emulating real-$E$ observables and near-axis resonances but can make isolating possible broad-resonance poles difficult.
+This is not problematic for emulating real-$E$ observables and near-axis resonances but can make it difficult to separate possible broad resonancs from the nearby states that supposedly represent BCT (s).
 This phenomenon needs further studies.
-Figure 4: The nonBorn term of the particle-dimer scattering $T$-matrix are emulated by two emulators with different $\mathrm{Im}\,E^{tr}_{\alpha}$.
+Figure 4: The non-Born term of the particle-dimer scattering $T$-matrix are emulated by two emulators with different $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$.
 For each curve, $\lambda$ and thus $B_{2}$ are fixed (see the legend), but a sample of $\lambda_{4}$ values is chosen to test the emulators.
 The means of the sample of relative errors are plotted.
-To emulate particle-dimer scattering amplitudes, we must vary the source parameters, including the interaction couplings and the scattering energy $E_{rel}$ between the particle and dimer, in addition to those in $H$.
-For on-shell scatterings, as computed in the emulation stage, the real-valued $E_{rel}$ is related to $E$ via $E_{rel}-B_{2}=E$.
+To compute and emulate particle-dimer scattering amplitudes, we must vary the source parameters, including the interaction couplings and the scattering energy $E_{\mathrm{rel}}$ between the particle and dimer, in addition to those in $H$.
+For on-shell scatterings, as computed in the emulation stage, the real-valued $E_{\mathrm{rel}}$ is related to $E$ via $E_{\mathrm{rel}}-B_{2}=E$.
 However, the training calculations treat them as separate emulation variables.
-Thus, we sample $N_{b}=128$ training points using LHS in the $\mathrm{Re}\,E$-$E_{rel}$-$\lambda$-$\lambda_{4}$ space, with $E_{rel}\in[0,60]$ MeV, $\lambda$ in a range in which $B_{2}\in[2,10]$ MeV, and $\mathrm{Re}\,E_{\alpha}^{tr}$ and $\lambda_{4}$ in the same ranges as for Fig.
+Thus, we sample $N_{b}=128$ training points using LHS in the $(\mathrm{Re}\,E,E_{\mathrm{rel}},\lambda,\lambda_{4})$ space, with $E_{\mathrm{rel}}\in[0,60]$ MeV, $\lambda$ in a range in which $B_{2}\in[2,10]$ MeV, and $\mathrm{Re}\,E_{\alpha}^{\mathrm{tr}}$ and $\lambda_{4}$ in the same ranges as for Fig.
 3.
-Two different emulators with $\mathrm{Im}\,E^{tr}_{\alpha}=3$ or $10$ are trained.
-They are checked against the exact on-shell $T$-matrices at a sample of 1000 points in the $E_{rel}$-$\lambda$-$\lambda_{4}$ space.
-Figure 4 demonstrates emulation performance by showing the relative errors of emulating the $T$-matrix’s nonBorn term versus $E_{rel}$.
-Each curve represents the mean of the relative errors from a sub-group of testing points with the same $B_{2}$ (see the legend) but different $\lambda_{4}$.
-The errors increase towards the particle-dimer threshold, similar to the behaviors of the rational approximations of functions near the branch points.
-Based on the comparisons, reducing $\mathrm{Im}\,E^{tr}_{\alpha}$ systemically improves the emulation, particularly around the branch points.
+Two different emulators with $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}=3$ and $10$ are trained.
+They are checked against the exact on-shell $T$-matrix non-Born terms at a sample of 1000 points in the $(E_{\mathrm{rel}},\lambda,\lambda_{4})$ space.
+In Fig.
+4, each curve represents the mean of the relative emulation errors from a sub-group of testing points with the same $B_{2}$ but different $\lambda_{4}$.
+The tiny errors at energies above thresholds suggest robust emulation performances there.
+Meanwhile, the (relative) errors increase towards the particle-dimer threshold, similar to the behaviors of the rational approximations of functions near their branch points.
+However, reducing $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ (from $10$ to $3$ MeV here) improves the emulations around thresholds systemically..
 Discussion.
-Our method has two components: the training calculations performed with complex $E_{\alpha}^{tr}$ and real $\bm{\theta}_{\alpha}^{tr}$, and the emulations that extrapolate the complex-$E$ training results to the other parts of the complex plane and interpolate (or extrapolate) in $\bm{\theta}$.
-The computing costs for emulation are low, but those of the training calculations increase with system size.
-Still, the complex-$E$ solutions have been computed in the continuum studies Carbonell et al.
-(2014) employing the CE Schlessinger and Schwartz (1966); Schlessinger (1968a, b); McDonald and Nuttall (1969); Uzu et al.
-(2003); Deltuva and Fonseca (2012, 2013a, 2013b, 2014) and LIT methods Efros (1985); Efros et al.
-(1994, 2007); Orlandini et al.
-(2014); Sobczyk et al.
-(2021, 2024); Bonaiti et al.
-(2024) for few- and many-body systems.
-Our method could thus apply to these systems as well.
-To infer the real-$E$ results from the complex-$E$ solutions, the CE method uses rational-approximation based extrapolants Schlessinger and Schwartz (1966); Schlessinger (1968a, b); McDonald and Nuttall (1969); Uzu et al.
-(2003); Deltuva and Fonseca (2012, 2013a, 2013b, 2014) while LIT fits real-$E$ observables via integral transforms Efros (1985); Efros et al.
-(1994, 2007); Orlandini et al.
-(2014); Sobczyk et al.
-(2021, 2024); Bonaiti et al.
-(2024).
-In our method, the RBM emulator is used to realize this connection.
-It creates an NHQM approach that computes not only observables but also spectra in which bound and resonance states are separated from the BCT states.
-All three methods discussed here thus belong to the type-II methods mentioned earlier—an insight of this work.
-Importantly, our method enables emulating the CE and LIT calculations.
-It provides fast interpolations of $\mathcal{A}$ at general $\bm{\theta}$ and $E$, given that $E$ is close to $E_{\alpha}^{tr}$.
-The complex-$E$ results at a given $\bm{\theta}$ can then be used as the inputs for the complex-$E$-to-real-$E$ procedures in CE and LIT.
-We thus create fast interpolations of the real-$E$ observables in the $\bm{\theta}$ space.
+Our emulators are trained using the calculations performed with complex $E_{\alpha}^{\mathrm{tr}}$ and real $\bm{\theta}_{\alpha}^{\mathrm{tr}}$.
+The cost of individual training calculation increases with system size, but its feasibility for few- and many-body systems have been demonstrated in previous studies [68] employing the CE [70, 71, 72, 73, 74, 75, 76, 77, 78]  [^6] Here, training points typically differ only by their $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ values.
+and LIT methods [2, 79, 3, 69, 80, 81, 82].
+These works underscore this powerful strategy that exploits advanced bound-state solvers to access continuum physics.
+The choice then becomes the method for inferring the real-$E$ results from the complex-$E$ solutions.
+To do so, the CE method uses continued-fraction based extrapolants [70, 71, 72, 73, 74, 75, 76, 77, 78] while LIT fits real-$E$ observables via integral transforms [2, 79, 3, 69, 80, 81, 82].
+In contrast, our method uses the RBM emulator to realize this connection, which keeps the number of training calculations minimal and offers extreme interpolating and extrapolating speed.
+Interestingly, this work creates an NHQM approach that computes both observables and spectra and obtains an insight that the complex-energy based methods belong to the type-II category mentioned earlier.
+Importantly, our $\bm{\theta}$-emulation, indispensable for exploring parameter space, can be adapted to work with the CE and LIT methods.
+It provides fast interpolations of $\mathcal{A}$ at general $\bm{\theta}$ and complex energies (close to $E_{\alpha}^{\mathrm{tr}}$), which can be used as the inputs for the complex-$E$-to-real-$E$ procedures in CE and LIT, creating fast $\bm{\theta}$-emulations of their continuum calculations.
 See Ref.
-Zhang (2024) for details.
-This work not only adds a new tool for studying continuum physics but also expands the MOR literature by applying the RBM method to project an operator with a continuum spectrum into a finite matrix; previous MOR literature Antoulas (2005) mainly focuses on dimension reduction of large matrices.
-Our study also suggests a connection between this RBM projection and (near-)optimal rational approximation of functions with branch points Nakatsukasa et al.
-(2018); Trefethen et al.
-(2021); Trefethen (2023); Nakatsukasa et al.
-(2023).
+[4] for details.
+Moreover, this work expands the MOR literature by applying the RBM method to project an operator with a continuous spectrum into a finite matrix; previous MOR literature [89] mainly focuses on dimension reduction of large matrices.
+Our study also shows great similarities between the complex-$E$ emulation and the ORA for functions with branch points [84, 85, 86, 87].
 Further studies.
-We derive essential understandings of the new method based on the training calculations with tiny errors (on the order of $10^{-12}$).
-Meanwhile, the numeric analytical continuation in $E$ is known to be sensitive to the errors in the data to be extrapolated Trefethen (2023).
-Such sensitivity will be addressed later.
-It will require the development of regularization methods, perhaps adapted to the specific many-body methods employed to perform the training calculations.
-For this purpose, the general behavior of the emulations exposed in this work, such as the distribution of eigenvalues, provides valuable prior knowledge for developing the regularization methods Hicks and Lee (2022).
+One direction is to implement greedy algorithms [46, 66] for selecting training points (including with random $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$), as suggested by the redundancy ($N_{\mathrm{eff}}\ll N_{b}$) observed above.
+We then obtain optimal bases with minimal training costs.
+The algorithms would also provide a more sophisticated regularization than the current SVD truncation, further stabilizing the analytical continuation in the complex-$E$ plane as seen in the context of ORA [84].
+The continuation, in particular near thresholds, could be also enhanced by enforcing general analytical behaviors in $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ within, e.g., the framework of Bayesian statistics [90].
+Lastly, we can extend the framework to treat systems with strong Coulomb effects by employing Coulomb wave functions [3].
 Summary.
-We apply the RBM method to emulate inhomogenous Schrödinger equations in a combined space of the complex $E$ plane and the parameters inside the Hamiltonian and the sources.
+We apply the RBM method to emulate inhomogeneous Schrödinger equations in a combined space of the complex-$E$ plane and other parameters in the equations.
 It creates an NHQM method for extracting and emulating continuum states based on bound-state-like calculations.
@@
 Good emulation performances are demonstrated in two and three-body systems as proof of principles.
-Moreover, the method can help emulate existing continuum calculations, such as those based on CE and LIT methods.
-Finally, the potential connection between the RBM’s effectiveness in mapping out the solution manifold and the optimality of the rational approximation should be further explored.
+The method can also help emulate existing continuum calculations, such as those based on CE and LIT methods.
+Finally, the potential connection between our emulation and ORA should be further explored.
 Acknowledgements.
 Discussions with Dean Lee, Dick Furnstahl, Nobuo Hinohara, Chong Qi, Simin Wang, Alex Gnech, Bijaya Acharya, and Ante Ravlic are greatly appreciated.
-I also thank Dick for careful reading of the manuscript.
+I also thank Dick Furnstahl for careful reading of the manuscript.
 This material is based upon work supported by the U.S.
@@
 DE-SC0013617 and under the STREAMLINE Collaboration Award No.
-DE-SC0024586.
+DE-SC0024586 (Michigan State University).